OpenStudy (anonymous):

Solve the following system and show all of your work. x − 2y = 14 x + 3y = 9

OpenStudy (anonymous):

subtract equation 1 from 2 we get 5y = -5 then put the value of y in either equation

OpenStudy (anonymous):

Hint: Substitution method. Solve for x in the first equation: x = 2y + 14 ^Substitute this in for the value of x in the second equation. x + 3y = 9 (2y + 14) + 3y = 9 Combine like terms and solve for y. Then substitute the value of y in either equation to find x.